### v3

# The Sunburst Arc with JWST. IV. The importance of interaction, turbulence, and feedback for Lyman-continuum escape

Abstract
At present, the best opportunity for detailed Lyman Continuum escape studies is in gravitationally lensed galaxies at . Only one such galaxy currently exists in the literature with sufficient magnification and spatial resolution: The Sunburst Arc at redshift . Here, we present rest-frame optical JWST NIRSpec integral field observations of the Sunburst Arc that cover a large fraction of the source plane. From this dataset, we generate precise maps of ISM kinematics, dust geometry, ionization, and chemical enrichment. We extract a stacked spectrum of five gravitationally lensed images of the Lyman-Continuum leaking star cluster, as well as a weighted, integrated spectrum of most of the galaxy, enabling a direct comparison to other LyC leakers in the literature. We find that the galaxy rotates but also shows strong, possibly dominant, signatures of turbulence, which are indicative of recent or ongoing major interaction. The cluster that leaks ionizing photons shows little difference from the rest of the galaxy in terms of kinematics or dust coverage, but dramatically elevated ionization, indicating that photoionization is the predominant mechanism that creates paths for LyC escape. We propose that tidal stripping of H I gas due to an interaction may have removed a large portion of the neutral ISM around the LyC emitting cluster, making it easier for the cluster to completely ionize the rest.
show]trive@astro.su.se
1 Introduction
A few hundred million years after the Big Bang, ionizing photons, termed Lyman-Continuum (LyC) radiation, from the first stars began to ionize the then-neutral IGM (e.g., hashimoto2018nature; haardt2012; fauchergiguere2020) in what is called the Epoch of Reionization (EoR). Passageways must thus have been present through the neutral ISM of these galaxies, which allowed LyC to leak out into the surrounding IGM, but it is not yet clear how these passageways were carved. In order to have reionized the universe, an estimated fraction of 10–20% of ionizing photons must have escaped their galaxies (robertson2015; naidu2020a), with some studies suggesting it might be more (becker2021; davies2021). The ionizing escape fraction cannot be directly measured in the EoR, because the IGM is opaque to ionizing photons at redshifts (inoue2014). All studies of escaping LyC must use lower-redshift galaxies to understand the underlying mechanisms. Such Lyman-Continuum Emitting galaxies (LCE) are rare in the local Universe, but become more common at higher redshifts, as both cosmic star formation activity and ionizing background grow larger. Since the first detection of a Lyman-Continum Emitting galaxy (LCE; bergvall2006), a small number of LCEs have been detected in the local Universe (see e.g. bergvall2006; leitet2011; puschnig2017; ostlin2021; komarova2024; leitherer2016); and at slightly higher redshifts, a considerable number of Green Pea (cardamone2009) and similar galaxies have found to be LCEs (e.g. borthakur2014; jaskot2019; wang2021; malkan2021b; flury2022a). At redshifts , about LCEs have been identified, depending on delineation criteria for candidates (vanzella2012; mostardi2015; debarros2016; shapley2016; vanzella2016; bian2017; chisholm2018a; fletcher2018; steidel2018; vanzella2018; riverathorsen2019; riverathorsen2022; saha2020; marqueschaves2021; marqueschaves2022; ji2020; flury2022a; pahl2021; saxena2022; kerutt2024).
Since it is not possible to directly measure galaxy LyC emission during the EoR, much work has been put into the identification and characterization of secondary observables which correlate with LyC escape, and the dual, intertwined questions of a) how they statistically scale with , and b) what they reveal about the mechanisms that carve out the paths of escape for the ionizing radiation. Very simplified, these mechanisms fall into the families of radiative versus mechanical feedback, which give rise to different escape scenarios. In a spherically symmetric, dust-free toy model of a young star cluster enveloped in an isotropic, neutral ISM, photoionization over time leads to a falling neutral fraction of the ISM in a growing Strömgren sphere which may eventually grow larger than the boundary of the ISM. This scenario is often called the “ionization bounded” scenario, where there is still an intact neutral envelope present, and the “density bounded scenario” when the Strömgren sphere has grown larger than the gas envelope (see e.g. zackrisson2013, for an overview). In such a scenario, depends solely on the column density of the residual neutral gas fraction in the ISM.
In contrast, radiative pressure and mechanical feedback such as stellar winds or momentum deposited from Supernovae, can result in large-scale ISM outflows, in which Rayleigh-Taylor instabilities will result in a breakup of this expanding bubble into a clumpy, highly anisotropic, expanding medium, in which some lines of sight will be transparent to ionizing photons, while others are optically thick. This is often referred to as the “perforated ionization bounded scenario” or the “picket fence scenario” (e.g. heckman2011; zackrisson2013; james2014; jaskot2014); in this limit, is regulated solely by the global or line-of-sight (LOS) covering fraction of the neutral gas. Real-life scenarios are of course much more complex, but the escape fraction can still be considered as regulated by the covering fraction of neutral gas clumps and the residual neutral column in the inter-clump medium.
A number of observables either correlate with, or directly encode, these two properties. One of the most direct measurements of the line-of-sight covering fraction is the equivalent width of H I absorption lines (heckman2011; jones2013; gazagnes2018; chisholm2018a), although for practical reasons, neutral metal lines are often used as proxies for these (e.g. riverathorsen2017; alexandroff2015; jones2013; borthakur2014; heckman2011). However, absorption lines are very observationally expensive, especially at high redshifts and thus often not practical for EoR studies. While neutral absorption lines directly trace the covering fraction, the properties of the Lyman- line directly traces the neutral column density in the inter-clump medium, although the strongly resonant nature of the line allows it to traverse open channels without any direct open lines of sight (e.g. gronke2016letter; dijkstra2016; dijkstrarev; verhamme2015; verhamme2008; verhamme2006). A number of line properties have been found to correlate with LyC escape, such as e.g. velocity offset of the red peak or separation of the dual peaks (verhamme2008; verhamme2015; jaskot2024a; jaskot2024b), or the steepness of the red peak (kakiichi2021). The presence of a narrow line component at systemic velocity has been predicted (behrens2014; verhamme2015; riverathorsen2017) and confirmed (riverathorsen2019) to indicate the presence of a line of sight of extremely low . However, due to the predominantly neutral IGM at the EoR, Ly emission cannot be depended on to trace ionizing escape at that time.
Other observables correlate either with high ionizing production rate, high escape fractions, or with physical conditions that bring about these conditions. In the latter category is the finding that LyC escape seems to correlate with broad emission line components, which in turn indicate the presence of large scale bulk outflows which can rupture and thin out the neutral ISM and create escape channels (see e.g. mainali2022, and references therein). However, in a study of low-redshift extreme emission line galaxies (EELGs), jaskot2019 found that outflows tended to be weaker at higher ionization, which they interpret as the onset of catastrophic cooling of neutral gas clumps which subsequently are photoionized by the still-proximate ionizing sources; marking the transition to a regime dominated by the density bounded scenario. LyC escape has also been found to correlate with star formation surface density (e.g. jaskot2024a, and references therein).
Besides properties internal to the source galaxy, the impact of the environment is a topic of active debate and research. Major mergers and IGM accretion events are predicted to spark strong star formation episodes, which in turn would increase the ionizing photon production, and feedback from this star formation could lead to increased escape fractions (e.g. kostyuk2024). Major interactions could also lead to strong turbulence, which in simulations have shown to be able to create ionizing escape paths through the ISM (e.g., kakiichi2019).
Observations have been ambiguous as to the importance of mergers. zhu2025 find from visual morphological classification of a sample of 23 LCEs in GOODS-South that 20 of them are mergers, a dramatic overrepresentation relative to the general merger fraction at (duan2025; puskas2025). In their simulated high-redshift IllustrisTNG galaxies, kostyuk2024 find that not only star formation induced by mergers, but also tidal stripping of the surrounding H I envelope plays a crucial role in facilitating LyC escape. Interestingly, this effect was observed directly by lereste2024, who used 21 cm interferometry to directly map H I in the local-Universe, interacting LCE Haro 11. These authors found that tidal stripping by the companion galaxy had offset large portions of the H I envelope away from the star forming knots in that galaxy, helping to clear out the ionizing escape paths.
Studying LCEs on the small physical scales required to map the ionizing escape paths is complicated by the fact that at redshifts low enough to allow the required spatial resolution, LyC emission is either completely absorbed by Galactic H I or, where the redshift is large enough to allow a window of escaping wavelengths, the localization of the LyC source is limited by the aperture of the Cosmic Origins Spectrograph; the only instrument which can effectively detect LyC at these redshifts. Samples at low redshifts redshifts are also limited by the small Cosmic volume that makes up the local Universe, as well as the low prevalence of Lyman Continuum Emitters (LCEs) in the present-day Universe.
At higher redshifts, where LyC has been shifted into the detectable wavelength ranges of HST imaging instruments, spatial detail is lost to distance. However, magnification by gravitational lensing can in fortuitous cases image galaxies at cosmological distances at a level of detail that is usually only seen in the local Universe. At the same time, at these distances, the ionizing radiation has been redshifted into ranges where it is comfortably observable in the WFC3 UVIS channel of HST, at physical de-magnified resolutions down to a few tens of parsecs or occasionally even better. As a result, such gravitationally lensed LCE galaxies currently present the best available opportunity for this kind of high-detail mapping of ionizing escape paths.
However, such lensed LCEs are rare. While searches are ongoing, at the time of writing, only one lensed LCE of sufficient brightness for detailed spatially and spectrally resolved studies is found in the literature; the Sunburst Arc (catalog ) (dahle2016; riverathorsen2019) at redshift . The Sunburst Arc was discovered during optical follow-up of Sunyaev-Zeldovich galaxy cluster candidates in the ESA Planck Survey (plancksz2014) by dahle2016, who found it to be the brightest known gravitationally lensed galaxy with an integrated magnitude of . Lensing by a redshift foreground cluster generates a total of 12 total or partial images (dahle2016; sharon2022; pignataro2021; diego2022; solhaug2025). sunburst2017 found that the galaxy displayed a unique, triple-peaked Ly emission profile, which is predicted to arise from Ly radiative transfer in an expanding shell of Hi perforated by a narrow channel devoid of Hi along our line of sight (behrens2014; verhamme2015). sunburst2017 estimated the neutral column density along this line of sight to be extremely low, , while one optical depth for LyC corresponds to ; orders of magnitude higher. The galaxy is dominated by a very young (3–5 Myr chisholm2019; vanzella2020; pascale2023; mestric2023; riverathorsen2024) and massive ( vanzella2020; riverathorsen2024) Lyman Continuum emitting (riverathorsen2019, from here referred to as "the LCE") star cluster. This cluster is dense and massive and contains a large population of Wolf-Rayet stars and Very Massive Stars (pascale2023; mestric2023; riverathorsen2024), and a strong enrichment in both doubly (pascale2023) and singly (welch2025) ionized Nitrogen, as has also been observed in a number of high-redshift galaxies with JWST (e.g. ji2025arxiv, and references therein). The LCE cluster has a very steep rest-frame UV slope of (kim2023), and mainali2022 found strong broad emission line components in ground-based, rest-frame optical spectroscopy, which those authors interpreted as the presence of strong bulk outflows driven by stellar winds from the young stellar population of the cluster. Based on this, mainali2022 concluded that mechanical feedback was the main driver of ionizing escape in the Sunburst Arc, presumably by driving the expansion and rupturing the surrounding neutral ISM. The present NIRSpec IFU observations provide an opportunity to spatially map the kinematic properties which have previously only been attainable in integrated slit spectra through atmospheric distortion, among other things enabling us to test this scenario more thoroughly.
This paper assumes a standard -CDM cosmology with km s-1 Mpc-1, . Unless otherwise noted, all images are oriented with N up and E left. We use the abbreviations [O I] for [O I] 6300, [O III] for [O III] 5008, [O II] for [O II] 3727+29, [N II] for [N II] 6584, [Ne III] for [Ne III] 3869, [S II] for [S II] 6717+31.
2 Observations
The JWST (gardner2023) data analyzed in this paper have also been included in the previous papers in this series; riverathorsen2024 , (riverathorsen2024), choe2025(choe2025), and welch2025. The observations were obtained with the NIRSpec instrument (boeker2023) in the IFS mode with the G140H and G235H gratings, and and with the NIRCam instrument (rieke2023). These observations were obtained over the period from April 10, 2023, to April 4, 2023, as part of program GO-2555 (PI Rivera-Thorsen).
The full field was observed with NIRCam in the filters F115W, F150W, F200W, F277W, F356W, and F444W. Four pointings were observed with NIRSpec; three on-target pointings and one off-target for background subtraction. Each of the NIRSpec pointings were observed in both the F100L/G140H and F170L/G235H settings, together covering the entire rest-frame Optical range at .
Figure 1 shows an RGB composite of the NIRCam F444W, F200W, and F115W imaging observations of the N and NW arc segments. Approximate footprints of the NIRSpec IFU pointings are overlaid in cyan, magenta, and yellow.
The NIRSpec pointings were designed such that P1 covers the largest possible part of the galaxy in one pointing; P2 is centered around the peculiar clump in its center, denoted “Tr” in vanzella2020, or “Godzilla” by diego2022; and P3 is placed to cover two strongly magnified images of the LCE cluster, to optimize the amount of information we can get about this cluster and its closest surroundings. Since the lensing shear of P1 is approximately perpendicular to the lensing shear in P2 and P3 (according to the lensing model of sharon2022), this choice of pointings optimizes the amount of spatial information that can be recovered.
Fig. 2 shows an artistic approximation of the de-lensed Sunburst Galaxy based on HST imaging, identical to the left panel of Fig. 11 in sharon2022, except here we have overlaid the approximate de-lensed IFU footprints; the direction in which each distorted square is shortest is the direction of its maximal spatial resolution. The figure consists of an image of the full, de-lensed arc as seen in the W and SE arc segments in sharon2022, onto which are superimposed by hand the more precise positions, colors, and brightnesses of the better resolved clumps found in the N and NW arc segments. The redder part in the N is a likely interacting companion galaxy, while the blue-green blurry parts in the center are likely to be a tidal bridge or similar feature.
NIRSpec Integral Field Unit (IFU) data were reduced following the methods presented in rigby23_overview. We used the JWST data reduction pipeline version 1.11.4 (bushouse23_pipeine1p11p4), with the calibration reference files from context pmap_1105. As discussed in rigby23_overview, the electronics of the NIRSpec detectors undergo small-scale fluctuations in temperature, which create pattern noise. We correct this detector pattern noise using the NSClean software (rauscher23_nsclean).
The raw NIRSpec data contains a large number of outlier pixels due to cosmic rays. The standard pipeline includes an outlier rejection step; however, this step does not adequately remove outliers from our final data products. We therefore use the baryon-sweep code (baryonsweep) to remove remaining outliers and artifacts. hutchison23_sigmaclip found that a combination of the standard pipeline outlier rejection and the baryon-sweep method produces the cleanest final data products.
To take full advantage of the magnifying effects of gravitational lensing, we produce two complementary data cubes, one with the standard spaxel resolution, and another with a finer spaxels. We use the Spec3 pipeline’s cube_build to produce these cubes, by setting the scalexy parameter.
Single-spaxel spectra of point-like sources from either of JWST’s integral field units (MIRI MRS and the NIRspec IFU) are susceptible to a “wiggle” artifact showing as faint, very broad waves in the continuum , which arises in each case because the detector does not critically sample the point spread function (law23_3ddrizzle). These wiggles are much broader than emission lines and easily removed in continuum subtraction, but they do give rise to these Moiré-like patterns in ratio maps. For the NIRSpec IFU, drizzling this undersampled data to even smaller output spaxels exaggerates the artifact (when using a standard 4-point dither as was done in this program). As noted by law23_3ddrizzle, using a larger extraction aperture around point sources mitigates this artifact, even in the NIRSpec data cubes. We find that the benefit from improved spatial resolution for extended sources outweighs the additional artifacts seen in point-like sources within this dataset. Hence, we have opted to analyze the 0.″05 data cubes.
3 Methods
3.1 Correction of Uncertainties
Fitting of emission lines in a spectrum requires a good characterization of the uncertainty spectrum. Underestimated uncertainties will lead to disproportionately strong weighting of the wings of strong emission lines, which in turn will lead to over-estimation of line widths, under-estimation of peak intensity, and poor recovery of the integrated line flux. We find that the data products produced by the pipeline have uncertainties under-estimated by a factor of for some spatial regions. Here we describe this measurement and how we rescale the estimated uncertainties.
We estimated the statistical uncertainty in every wavelength bin in the Level 3 datacubes as follows. We first manually constructed masks covering regions of empty sky. These masks were placed to minimize line emission from the source galaxy, but were also kept well clear of edges of each pointing, since these regions are noisier and not representative of the noise properties of the spaxels covering the target. The masks are shown in Figure 13.
For each wavelength bin, we then took the standard deviation of the data within this mask, after first sigma-clipping the data by to avoid artifacts like hot pixels and cosmic rays from skewing the distribution. We smoothed the resulting error spectrum by a running median kernel to remove any remaining individual bins of inflated error.
For each spaxel, we then assumed a new uncertainty spectrum, as the greater of the sky-region–derived uncertainty and the pipeline-provided spectrum. This ensured that individual wavelength bins and spaxels with higher uncertainty still were treated correctly, while the floor of the uncertainty is set by the measured standard deviation in sourceless pixels.
3.2 Stacked spectrum of the LyC leaking knot
Of the twelve gravitationally lensed images of the LCE cluster, five are fully captured by the NIRSpec IFU pointings. Since these are all physically the same object, we combined their spectra by extracting each in a pixels aperture, normalized each spectrum by its median value to standardize their differently magnified fluxes, before stacking them by taking a flux-weighted average. All quantities derived from this spectrum are thus given in relative or units.
Figure 3 shows the full stacked, rest-frame blue and optical spectrum of the LCE cluster. The spectrum contains a large number of emission lines; we have identified 59 nebular lines, of which 30 are Hydrogen recombination lines (including about 20 Balmer lines and 10 Paschen lines). The spectra also feature a set of broad, blended, stellar Wolf-Rayet line complexes that were discussed in riverathorsen2024. The Figures in Appendix B identify all these features. To enable conversion to physical units, we report the measured and the demagnifed flux of H for the image of the LCE that appears in image 4 of the lensed galaxy: the measured flux is , and the demagnfied flux is , assuming the calculated magnfication of from K. Sharon et al. (2022).
3.3 CubeFitter.jl: A new software package for fitting line fluxes and kinematics in spectral cubes
For spectral line modeling and generation of maps of line flux and kinematics, we used the Julia package CubeFitter.jl (riverathorsen2025cubefitter)111https://github.com/thriveth/CubeFitter.jl, a software package that quickly and flexibly models emission lines in spectroscopic datacubes. CubeFitter is written in the Julia programming language (bezanson2017julia), to strike a balance between computation speed and simple, flexible development. Julia has proven for our tasks to be faster than Python, while taking up a number of lines of code and a level of coding complexity similar to Python code of equal functionality. CubeFitter relies on a number of existing Julia packages for its core functionality, most importantly the general curve fitting package GModelFit.jl222https://gcalderone.github.io/GModelFit.jl for model building infrastructure and fitting functionality, Measurements.jl333https://juliaphysics.github.io/Measurements.jl for propagation of uncertainties, and optionally VoronoiBinning.jl444https://github.com/Michael-Reefe/VoronoiBinning.jl, a Julia implementation of the original Vorbin package (capellari2003).
The CubeFitter.jl package loads a FITS datacube into a data structure with relevant, partially user-provided, metadata such as instrument resolving power, data units, and more. The program ships with a list of emission lines, mainly taken from Drew Chojnowski’s online line list555http://astronomy.nmsu.edu/drewski/tableofemissionlines.html, in turn compiled mainly from data from NIST (nist_asd). A few additional lines are taken from other places, see the package documentation. Alternatively, the user can pass their own line list to the code. The package incorporates the tables of spectral resolution for NIRSpec666available for download at https://jwst-docs.stsci.edu/jwst-near-infrared-spectrograph/nirspec-instrumentation/nirspec-dispersers-and-filters#gsc.tab=0. These are pre-launch estimates; although in-orbit measurements have recently been made (shajib2025arxiv), they have not yet yielded data products suitable for implementation in code like this..
An optional function allows the user to subtract the continuum under the emission lines. The function, cont_subt, estimates the continuum by masking out a wavelength range around a set of the strongest emission lines, sigma clipping the remaining data bins to 3 , and taking a running median within a broad kernel (51 pixels by default, but settable through an option argument to the function). This running median is then interpolated into the masked pixels and subtracted from the original spectrum, and the procedure is repeated for better masking and sigma clipping. Caution should be given to pick a median window width suitable for the data; specifically, it will subtract all features with a width this window.
CubeFitter can either extract individual spaxels, or co-add arbitrary rectangular cut-outs or image segments into one-dimensional spectra. When co-adding spectra, the uncertainties are expected to be random and independent in neighboring spaxels and wavelength bins, such that they can be propagated by addition in quadrature in the usual way.
Emission lines in an extracted spectrum are fitted simultaneously with Gaussian profiles, assuming that all lines follow the same kinematic structure with a shared redshift and velocity width, leaving line flux as the free parameter for each line. By default, the code attempts to fit all lines in the line list. Optionally, a custom subset of these can be passed, which enables fitting of multiple subsets or individual lines to different kinematic parameters. The code can also fit any line or set of lines with locked kinematic properties, or fit the kinematic properties from a subset of lines and use that fit as a kinematic template for the rest. These features allow for very flexible modeling strategies.
Data regions with undetected lines contribute no signal to the fit, increase computation time, and unnecessarily increase uncertainties of the resulting fit parameters. To avoid this, the code performs a quick, numerical estimate of the signal-to-noise ratio of each line in each spaxel, and discards any lines below a (settable) S/N threshold, inserting a NaN value in the resulting flux maps for the given line. The code has rudimentary support for modeling two kinematic components in each line, but at the time of writing this, it is not yet possible to combine this with the use of a custom selection of lines to compute the kinematics.
3.4 Emission line maps
3.4.1 Line fitting strategy
For each spaxel and filter/grating setting, we modeled all emission lines as single Gaussian profiles, with a shared redshift and velocity width. We constructed this model as follows:
We picked a number of the strongest lines on which to base the kinematics. In the case of F100LP/G140H, we used [O II], H, and [O III]; and for F170LP/G235H we used [O III], H, and [N II] (blended features must be fitted together to get correct fluxes). These were then used to build one single model, which was fit to the observed data. We have not attempted to account for the possible presence of multiple gas components in each line, and we have not attempted to account for possible kinematic differences between various ionization zones. A more detailed and in-depth kinematic modeling of the gas is deferred to a follow-up paper.
Before fitting the spectrum of each spaxel, we removed data chunks with negligible line flux, to avoid empty chuncks of data degrading the signal-to-noise of the resulting fits. To this end, we numerically estimated the integrated flux and standard error within a km s-1 window around each line. If this yielded a S/N < 0.5, the line was excluded from modeling in this spaxel.
After simultaneously fitting strong and blended lines together in one model, we cycled through a number of weaker lines and modeled each individually, locking each to the kinematic properties found from the strong lines and allowing only the line flux to vary.
3.4.2 Masking
Because of the strongly elongated shape of the arc, a large number of spaxels in each datacube do not contain any emission from the source galaxy. These spaxels may however still spuriously give rise to fluxes above the S/N limit set during line fitting. We have thus constructed masks to remove pixels based on the following criteria:
1.
Since the kinematics in most pixels is dominated by the [O III] 4960,5008 Å doublet, we produced a stacked [O III] S/N map by assuming a fixed line ratio of (vanhoof2018) and, once this fixed line ratio was corrected for, taking the mean of the four line maps (each doublet member in each of the two settings). Uncertainties were propagated in the standard way. We then created a S/N map, and excluded all spaxels with for this stacked [O III] map.
2.
Because there was still a significant number of spurious spaxels present in the resulting mask, we produced a polygon mask by hand, using the S/N mask as a guide. This mask is shown in Figure 14.
These masks have been applied to all maps shown subsequently, as well as all statistical treatment of source spaxels, unless otherwise stated. For many lines and line-derived property maps, additional spaxels outside these masks also contain missing values due to lower S/N ratio in one or more of the involved lines, but spaxels within the masks have been removed everywhere.
3.4.3 Combined maps
The F100LP/G140H and F170LP/G235H observations resulted in separate datacubes which we did not attempt to combine directly, but produced separate line and kinematic maps of each cube. This resulted in duplicates of the maps of kinematic properties, as well as flux maps of the [O III] 4960,5008 Å lines which are included in both grating/filter settings. We combined these maps by a simple average of all spaxels present in both frames of the relevant quantity, with standard errors propagated by summation in quadrature. Whenever a pixel had a missing value in one of the maps, this pixel in the combined map would be set equal to the one existing value.
3.5 Dust attenuation maps
We have accounted for dust attenuation in the galaxy using a standard starburst attenuation law (calzetti2000). We have computed using the Balmer decrement of H/H following standard methods as outlined in e.g. dominguez2013, and using intrinsic line ratios from osterbrock assuming and . In the very high density region surrounding the “Godzilla” region (see choe2025), the emission from partially ionized regions (dai2024arxiv) or broad H emission from the power-law winds of the extremely lensed Godzilla object may lead to an elevated H/H line ratio and thus an overestimation of ; a more thorough treatment of dust attenuation in this region was done by choe2025, who found a strongly elevated from H in this clump, while a selection of other Balmer and Paschen lines unaffected by this effect yielded a value of only half that. We therefore advise caution when interpreting the values found from H near the Godzilla region. However, the rest of the areas covered by the IFU are more typical for warm starburst ISM, and thus a H based determination of is appropriate for this region.
From these dust reddening maps, we then produced maps of dereddened line flux of all fitted emission line maps. However, the NIRSpec IFU pixel size undersamples the PSF, even in the pseudospaxels resulting from drizzling the 4-point dithering pattern used for the present observations. This may lead to a fringing-like stripe pattern when dividing line maps (see e.g. figures in Sect. 3.6. We have therefore used the dereddened maps only when it was necessary, and we caution the reader that this residual stripe pattern is not physical.
3.6 Determining ionization and abundances
In addition to dust and kinematic properties, we have computed spatially resolved maps of commonly-used diagnostics of ionization conditions and chemical abundance. These are the line ratios , as well as the ionization diagnostic , and the diagnostic line ratio . From S2, we also computed a spatially resolved map of the sulfur deficiency described in Sect. 3.7.3 below. O3 and are shown in figures in Sect.3.5; the rest of these maps are not shown, but were included in the analysis of the next section.
3.7 Reconstructing global ISM properties of the source galaxy
The advantage of spatial resolution motivates this paper. However, in order to contextualize the results with samples of unlensed galaxies, we need to compute its integrated or global properties, as it would appear if not lensed and observed at ordinary spatial resolution.
3.7.1 Masking and de-magnification map
To emulate properties of the galaxy as they would have been observed in a slit spectrum without gravitational lensing, we extracted the global spectroscopic properties of the galaxy by the following method. First, we constructed a mask which contained only the one lensed image that covers the largest part of the source galaxy. The region, which is found in P1, is shown in Figure 4. This galaxy image is bounded in both directions along the arc by crossing critical curves, and in the off-arc direction predominantly by the [O III] S/N based mask from subsubsection 3.4.2.
The location of the critical curves is derived from the lensing model of sharon2022. To compensate for differential magnification within the mask, we used magnification maps derived from the same work. We used the Python package Reproject to project the magnification maps onto the WCS of the drizzled NIRSpec cube of P1. We then applied the combined S/N and critical curve mask to the magnification map as well. The inverse of the resulting masked map can then be used to compensate for differential magnification. We note that the exact delineation of the mask along the critical curve is of limited importance: Magnification is high in the vicinity of the critical curve, and since we weigh the data by the inverse magnification, a few pixels too many or too few in this region will have very limited impact on the resulting values of integrated and weighted quantities.
A more thorough treatment of a magnification-weighted, integrated spectrum and comparisons to non-resolved starburst galaxies at similar redshifts is deferred to future work. Here, we have used it to derive global values of properties derived per-spaxel in this work; including redshift, ionization parameter, [S II] deficiency (see subsubsection 3.7.3), and more. An important caveat to these properties is that the regions included in the mask in pointing 1 are not a fully complete image of the galaxy. As is evident from Figure 2, the image shown in Figure 4 covers most of the galaxy, but not all. All global properties are given under the assumption that the included regions are an accurate representation of the global properties of the galaxy.
3.7.2 Global kinematic properties
A common means to characterize the kinematical properties of a galaxy is by comparing the shearing velocity to the intrinsic velocity dispersion (glazebrook2013; herenz2016; bik2022; herenz2025). is defined as the difference in centroid velocity between the spaxels with the most extreme values. is defined as the flux-weighted velocity dispersion averaged over all spaxels. Following herenz2016, but substituting [O III]5008 for H, these are written as:
Following herenz2016, we have used the 5th and 95th percentiles of velocities to avoid outliers from skewing the results, while still sampling the extreme ends of the velocity distribution. We opted to base on [O III]5008 rather than H, because in our dataset, the former both intrinsically has higher S/N than the latter, and because this line at the redshift of the Sunburst Arc falls in the overlapping range between G140H and G235H and thus was observed in both, effectively doubling the exposure time. When computing , we have not made any attempt to correct for beam smearing such as has been done in the ground based studies of e.g. glazebrook2013; herenz2016; bik2022; herenz2025. The effective spatial resolution of JWST/NIRSpec for this lensed target is so high that the problem is negligible. Additionally, the beam smearing correction in these works was motivated by a central spike in line width in those kinematics maps caused by the atmospheric blurring, leading to an overestimation of the brightness-weighted velocity dispersion in these works. We do not observe such an effect, and consequently do not expect the same skewing of .
The observed value of is the projection along the line of sight of the true, physical value, and needs to be corrected by a factor of , with being the inclination angle. Finding the inclination of the galaxy requires detailed 3D modeling, something which is outside the scope of the present work. However, we can apply some statistical and geometrical considerations. If we assume a disk-shaped galaxy at an inclination , we know that the median inclination angle in a large sample of randomly oriented disks is 60°. If we adopt this value, this gives a correction factor to of .
3.7.3 Ionization properties and diagnostic line ratios
Using the single-galaxy mask and inverse magnification weight map, we have computed global values of diagnostic line ratios and derived properties that are often used in the literature, so that the Sunburst Arc can be directly compared to observations of unlensed galaxies at similar redshifts. These also provided the opportunity to compare the properties derived from the stacked spectrum of the LCE, to a spectrum as it would more realistically look if unlensed and unresolved. This could tell us whether the Sunburst Arc is globally similar to the LCE galaxies found in LCE surveys (e.g., nakajima2020; cooke2014; fletcher2018; liu2023; wang2021; xu2023; flury2022a; flury2022b).
We calculated each of these diagnostics by extracting a inverse magnification weighted spectrum in the single-image mask discussed in Sect. 3.7.1. We then used these weighted, masked and integrated line fluxes to compute derived properties. In addition to the global kinematic properties and the spatially mapped diagnostics discussed in Sect. 3.7.3, we computed the sulphur deficiency as the difference between and the locus described in wang2021. In a similar manner, we computed global values of and derived properties.
4 Results
We first present spatially resolved maps, then present results based on the weighed and integrated maps described in subsubsection 3.7.1, as well as the stacked LCE spectra described in subsection 3.2.
4.1 Line flux maps
Figure 5 and Figure 6 show the emission intensity maps for the majority of the lines used in the diagnostic maps that follow. Each line map is colored to loosely correspond to the rest-frame wavelength of the line. The masks described in subsubsection 3.4.2 are applied here and in all subsequent maps, unless otherwise stated.
To optimize visual quality, we show the sum of both lines in the [N II] 6548,84, although only [N II] is included in further diagnostics.
4.2 Kinematics
Maps of centroid velocity relative to systemic () and the velocity width (FWHM) resulting from the fitting described above in Sect. 3.4.1 are shown in Figure 7, along with maps of stellar continuum shown in a separate column, as well as overlaid as contours on the kinematic property maps as an aid for visual orientation.
In Figure 7, the map of P1 (upper center) shows a clearly defined velocity gradient from E to W which, as seen in Figure 2 roughly corresponds to the same direction in the source plane. However, this apparent simplicity might be slightly misleading. The parts of the galaxy which are redshifted relative to systemic velocity (to the E, shown in orange tones) has a somewhat more complex kinemorphology. In addition, due to their brightness, the blueshifted parts are dominant in the integrated spectra from which the systemic velocity is derived. If we focus solely on the single, almost-complete image of the galaxy (see Figure 4), the blueshifted parts make up only a small part of this. Thus, if we had compared to a simple, unweighted average of the centroid velocities, the kinematic complexities in the regions of the galaxy not containing the LCE would likely have appeared more complex than they do in Figure 7.
The right column of Figure 7 shows the velocity-space FWHM. Of particular interest is the dramatic line broadening immediately to one side of the multiple images of the LCE but, interestingly, not actually coincident with it. Comparing to the middle column of Figure 7, we see that this region of high FWHM is coincident with a slight spike in outflow velocity. This offset kinematic spike might very well originate in a separate outflowing gas component which is automatically incorporated into our single-component model. In P2+3, this region appears particularly dramatic and elongated immediately NE of the LCE image 1.8 (in the terminology of sharon2022), immediately next to the object “Godzilla”. This region is subject to extreme magnification and shearing due to a low surface brightness lensing foreground galaxy (choe2025), but despite this extreme magnification and distortion, it is still another lensed image of the same region as in P1.
4.3 Dust and ionization
In Figure 8, we show spatially resolved maps of dust reddening and ionization diagnostics, again with the upper row showing the maps for P1, and the lower row showing them for P2+3; and with stellar continuum maps shown in the leftmost column.
In the center-left column, we show maps of derived as described in subsection 3.5 from the H/H ratio and a Calzetti attenuation law. The stripey patterns showing in these panels are an artifact of the undersampled PSF. We see that the dust attenuation is quite uniform across the arc, except in the “Godzilla” feature marked in Figure 1, and its much fainter counterimages in both P1 and the lower left of P2+3 (see choe2025, for a further discussion).
The third column shows maps of the ionization diagnostic [O III]/[O II]; the same artifacts are present here as in the maps of . The ionization is quite elevated around the images of the LCE and in the central part of P1, we also note the feature in very N end of P2, which is associated with a very faint star cluster, yet apparently displays an extreme ionization of .
Finally, the right panels show another ionization diagnostic, , which has been proposed as an improved tracer of LyC escape due to the lower ionization potential of [S II] compared to [O II], meaning that these lines more accurately trace the presence or absence of neutral gas, where O+ ions also are present in substantial amounts inside H II regions and the photodissociation regions at their boundaries. While the [S II] line may be fainter than [O II], the diagnostic is based entirely on lines very close in wavelength and thus independent of dust model. For the same reason, it these stripey artifacts are suppressed in this map, although still visible. It is to our knowledge the first time it has been possible to map both log [S II]/H and LyC emission in any galaxy and thus test its strength as a tracer on a local scale. We note that there is a strong similarity between the spatial distribution of high- and low-ionization regions by the two methods.
4.4 Nitrogen spike
Figure 9 shows a maps of N2 = [N II]/H. We note the conspicuous regions of elevated [N II]/H at locations next to, but not coincident with, the LCE cluster. In these regions, the elevated [N II] might be due to elevated N/H, or due to collisional excitation from shocks. Spatially, they closely match the regions of larger line FWHM and a slight rise in blueshift as seen in Figure 7; something which could support the collisional interpretation. The region is not coincident with any star clusters we have been able to identify in our NIRCam images or archival HST images; we speculate that it might be the result of an anisotropic outflow from the LCE cluster.
4.5 Global properties
In addition to the spatially resolved maps, we have computed values of some core diagnostics from the stacked LCE spectrum, as well as globally for the full, magnification-weighted galaxy as described in Sect. 3.7.1, allowing for a more apples-to-apples comparison of properties of the Sunburst Arc galaxy to other known LCE galaxies in the literature and a discussion of global vs. local properties in the galaxy and the efficiency of observational proxies for LyC escape. The properties and their measured values are tabulated in Table 1.
| Quantity | Galaxy | LCE |
| [km s] | — | |
| [km s] | — | |
| O32 | ||
| R3 () | ||
| O1 () | ||
| N2 () | ||
| S2 () | ||
5 Discussion
5.1 Kinematics
The upper center panel in Figure 7 clearly shows that the galaxy has a distinct, ordered velocity gradient. As this pointing (P1) has the most homogeneous magnification and least complex distortion pattern (see Figure 4) and thus the lensed image within this mask likely is quite similar in shape to the source gaalxy, this translates quite smoothly into a gradient in the direction from the lower right to the upper left corner of the de-lensed P1 footprint shown in cyan in Figure 2, the kinematic structure of a rotating disk. The picture is more complex in P2+3 (lower panels of Figure 7), but this is expected: The shearing direction in the lower part (P3) is at a high angle to that of P1, and in the upper parts, the lensing geometry is complex and not well solved. As discussed in Sect. 4.2, we have found a shearing velocity of about 109 km s-1, and a flux-weighted velocity dispersion .
If we adopt the median inclination angle reached above, value, this gives a correction factor to of , yielding a corrected value of . In (schreiber2020; bik2022), a value of is adopted as the demarkation above which they consider a galaxy to be rotation dominated. This value is, for the Sunburst Arc galaxy, reached at angles of (see section 4). In a random distribution, one finds about 30% of galaxies in this inclination angle range, meaning that the galaxy, assuming no knowledge of its orientation, has about a 70% probability of being dispersion rather than rotation dominated, and for the majority of the remaining 30% not very strongly rotation dominated.
This paints a picture of a galaxy which is rotating, but also undergoing significant turbulence. From the FWHM maps in Figure 7, the velocity disperson does not spatially correlate with strong star formation activity (apart from localized and directional outflows near the LCE, cf. section 4.2 and 4.4), suggesting that this strong dispersion, rather than stemming from star formation feedback, may be due to outside influences such as tidal disruptions by the Northern companion seen in Figure 2.
5.1.1 The origin of a broad emission component in H and [O III]
It has long been predicted that stellar winds and supernova feedback would play an important role in clearing the escape paths for Lyman-continuum, both in theoretical works (trebitsch2017; rosdahl2018; kakiichi2021) and in observations (heckman2011; chisholm2017; kim2020). amorin2024 have found a strong preponderance of broad emission lines in LyC-leaking Green Pea galaxies, although it has been difficult to establish any direct correlations between broad emission components and LyC escape fraction.
mainali2022 studied the rest-UV and -optical emission in the Sunburst Arc from the ground using the Magellan/MagE and Magellan/FIRE spectrographs, respectively. These authors did find a strong broad component in the emission features of H and [O III] in the LCE spectra, while spectra from non-leaking regions showed a considerably fainter broad feature. However, the broad feature in mainali2022 has FWHM km s-1 or km s-1 somewhere in between the broad- and medium-width features of (amorin2024). Compared to the ground-based spatially unresolved spectra of mainali2022, the current NIRSpec IFU observations enable spectra to be extracted from spatial regions much more selectively, with lower contamination from nearby regions.
riverathorsen2024; welch2025 extracted and stacked spectra from multiple lensed images of the LCE, allowing for unprecedented signal/noise ratio and absence of contamination from surrounding regions. Both these works modeled strong, rest-frame optical nebular emission lines using three Gaussian components, yielding FWHM values of km s-1; consistent with each other and broadly consistent with typical values of the narrow, medium and broad components found by amorin2024.
In this work, we have only used one Gaussian component to model the emission lines in each spaxel, so our measured velocities are not directly comparable to those of the above works; in particular, faint but broad components will have limited impact on the model line width and centroid velocity. On the other hand, single component fits can still robustly show spatial variation in kinematic properties, and a brightening of a broad component will manifest itself as a larger measured line width. Looking at the FWHM maps in the right panels of Figure 7, it is interesting to note that while we know from the stacked spectra that a broad component is present in the LCE cluster, it is not strong enough to lead to an elevation in the FWHM map; which is dominated by a component of km s-1 consistent with the surroundings.
The line is however substantially wider in a small region immediately next to the LCE as seen in the two bright spikes in P1, reaching an FWHM of km s-1. The same region appears twice in P2+3 in highly elongated form due to the strong lensing shear in these images. This region is also slightly blueshifted relative to the LCE () and shows an equally conspicuous spike in the [N II]/H ratio. It is not immediately clear what drives this apparent outflow, as it is spatially distinct from the LCE and there is no star cluster visible at its location. It is close enough to the LCE to have bled significantly into the MagE and FIRE slits used by mainali2022. We therefore conclude from our maps that this region, rather than the LCE itself, gives rise to the broad component in the LCE spectra in that paper. The ground-based spectra of mainali2022 lacked the spatial resolution to make this distinction. If correct, this interpretation of course raises the question of how often the observed broad components in LCEs from the literature originate from nearby regions, rather than from the isotropic stellar winds of the leaking clusters themselves, and in extension, whether a wider array of mechanisms are needed to explain the appearance and sustenance of ionizing escape channels.
5.2 Dust geometry
The leaking clump is known to have a bright, narrow Ly emission spike at systemic velocity (sunburst2017). This is indicative of have a line of sight towards us that is virtually free of neutral gas. However, the dust reddening as traced by the Balmer decrement is not significantly lower along this line of sight compared to others in the vicinity. This lack of a hole in the dust, together with the relatively smooth and undramatic behavior of the kinematic parameters around the LCE, indicates the absence of strong mechanical feedback. If mechanical feedback is not at work, this implies that photoionization is the main mechanism creating the escape path for Ly and LyC photons, at least in the immediate surroundings of the LCE. As discussed above in Sect. 5.1, the suggests that the galaxy is currently or recently interacting with anothet galaxy. As discussed below, an interaction could have helped strip the bulk of the surrounding H i envelope away from the region surrounding the LCE.
5.3 Ionization
As discussed in section 1, the diagnostics is a good indicator of overall ionization in a given region, but not a good predictor of LyC escape. An alternative, ionization-based tracer of LyC escape is the relative weakness of [S II], first suggested by alexandroff2015. Like the O32 diagnostic, it is a measure of ionization; but the ionization potential of neutral sulphur is as low as 10.36 eV, lower than the ionization potential of hydrogen, such that it is effectively shielded from ionization in neutral regions. [S II] lines are only emitted from only present at the outer boundary of H II regions. These boundaries disappear as a region becomes fully ionized, exactly in the conditions that allow LyC escape.
For the Sunburst Arc, we can compare the spatially resolved maps of log [S II]/H shown in the rightmost panels of Figure 8 to both the local values at the LCE cluster, and the global, integrated value as found in the previous section. From Figure 8, it is clear to see in the rightmost panels that the [S II]/H map has a strong, and highly localized, minimum at all the images of the LCE cluster. From Table 1, we see that the local value at the LCE is indeed , consistent with what is seen in the figure, while the global value is . These latter two values correspond to values of the quantity [S II] as defined in wang2021 of and , respectively. Comparing to the histograms and derived kernel density estimates in Fig. 2 in wang2021, we see that these values put the LCE into a range in which the LzLCS contains practically no non-LCEs, while the value from the integrated spectrum places it slightly above the average from SDSS. This value falls in a tricky region in the KDEs of wang2021; it is on one hand dominated by by non-leakers, but on the other hand it is also close to the peaks of the distributions both of strong LCEs and that of all LCEs in the LzLCS.
Thus, the LCE cluster alone has such strong [S II] that it could not be mistaken for a non-leaker. In contrast, the galaxy as a whole displays a [S II] level which is almost equally represenative for leakers and non-leakers. Thus, we see that while [S II] seems to be a very good local tracer of ionized channels, the diagnostic can still be diluted by surrounding regions, making a leaker look like a non-leaker, as also found in wang2019; wang2021.
The strong contrast in [S II] at the LCE compared to the surroundings also agrees very well with the predictions, based on which the method was initially proposed, and supports that the method is indeed well physically motivated.
For comparison, Figure 10 also shows the values of [S II] from the stacked archival JWST spectra of star forming galaxie from hayes2025. The stacks are grouped by EW([O III]) as a proxy of ionization. While the Sunburst Arc LCE stack sits somewhere between the fourth and fifth quintile of the high- ionization distribution, the global spectrum is more modest. Interestingly, though, even the galaxy as a whole has higher ionization and stronger sulfur deficit than the majority of observed galaxies in the high-{z}.
5.4 Global ionization properties
We also produced the classic Baldwin, Philips, Terlevich (BPT baldwin1981; kewley2001; kauffmann2003; kewley2006) diagram of the Sunburst Arc based on line ratios extracted both locally in the LCE cluster stack, and globally from the weighted map described in subsubsection 3.7.1, left panel. For comparison, we also included the LzLCS (wang2021; flury2022b), with their LCEs shown as filled and their non-LCEs as empty circles. The shaded zone in the upper left corner of the diagram is the erb2016 Extreme Emission Line Galaxy (EELG) classification region. Interestingly, while the Sunburst Arc on the diagnostic maps in e.g. Figure 8 can look like an overall quite modestly ionized galaxy, once the diagnostics properties get weighted by flux as they would in a spatially integrated spectrum, it is revealed as a quite extreme galaxy, even for its redshift, which is noteworthy considering the general breakdown of the correlations between just these these properties and LyC at .
In addition to the classical BPT diagram, we have also included the [O I] based diagram in right panel of Figure 11. This diagram shows that there is little sign of shock ionization of the ISM in the galaxy, at least not in the brightest regions in the LCE neighborhood; consistent with the a general absence of strong mechanical feedback in this region.
5.5 LyC escape scenario
sunburst2017 observed a triple-peaked Ly profile emanating from the Sunburst Arc, consisting of the classic double-peak profile observed many times in the literature, combined with a narrow peak with a peak velocity and line width almost identical to those of H, telling of a component undergoing no radiative transfer effects, escaping through a channel of less then one optical depth in Ly. Based on this, sunburst2017 predicted a scenario in which Ly escapes directly through a narrow perforation in an otherwise optically thick shroud of H i, in a manner reminiscent of the arc’s namesake meteorological phenomenon. According to our best models at the time, the total solid opening angle could be no more than than 2-5% of the total projected area surrounding the LCE cluster, with an extremely low neutral column density of . If the combined opening angle were larger, practically all Ly photons were predicted to escape through the open channel without the wavelength-space redistribution leading to the observed dual peaks surrounding the central peak. Such a channel would have to be created by a highly directional effect such as a jet from a previous accretion event; and it would have to be coincidentally aligned to a high degree with our line of sight. While such a scenario is certainly possible and might indeed have been the only plausible explanation; having to invoke this level of coincidental fine-tuning is uncomfortable, and furthermore, such a puncture by jet scenario would also have cleared away the majority of dust along the line of sight, but a significant dust screen was already then observed to be present in the nebular emission in sunburst2017, and since confirmed in models of the stellar population (chisholm2019; riverathorsen2019).
Recent modeling work by almadamonter2024 has shown that H i clouds are much more “sticky” to Ly photons than previously believed, in the sense that Ly photons have a substantially higher probability of getting trapped in the H i gas and undergoing significant spatial and frequency-space redistribution, than was believed at the time of sunburst2017; riverathorsen2019. Consequently, a much wider range of opening angle than previously believed can give rise to the characteristic Ly profile of the Sunburst Arc; perhaps angles as large as 30-40% of the total surrounding solid angle. In this case, there is no longer a need to invoke a purely serendipetous, á priori very unlikely alignment of a jet or similar, highly directional, mechanism with our line of sight, and with this proposition also disappears the problem of explaining how such a case of mechanical feedback would have extremely efficiently cleared a pathway of neutral gas, yet left a seemingly unaltered dust screen behind. With the new anisotropy characteristics allowed by almadamonter2024, we once again find ourselves in territory in which photoionization combined with an asymmetric H i geometry could clear away the H i in a cone-like configuration, while leaving the dust screen largely unchanged, as well as leaving behind a majority of projected surrounding area still containing enough H i to make it optically thick to Ly, thus generating the classic double-peaked Ly profile surrounding the central peak.
We propose that this asymmetric H I geometry is provided by the same kind of tidal stripping from a major interaction as found in Haro 11, discussed in section 1. Comparing Figure 2 to Fig. 1(a) in lereste2024, the Sunburst source galaxy has a geometry consistent with a similar scenario. In the upper (N) end is the interacting companion, connected to the main galaxy with a bridge-like structure in the central part of the figure. The LCE cluster is located in the bottom (S) end of the figure, farthest away from the companion in a region that would be stripped of much of its H i in a scenario similar to Haro 11. We therefore propose that tidal stripping, while not shown to actually take place, could plausibly be occurring also in the Sunburst galaxy and could well be at least in part responsible for clearing the paths for the escaping ionizing radiation. The high ionizing photon production rate in the LCE cluster has helped ionize the remaining neutral ISM with extreme efficiency, given the extremely low neutral column density inferred from the Ly profile. The seemingly even dust cover suggests that photoionization, rather than mechanical feedback or radiative pressure, has played the main role in clearing the ionizing escape path.
A conceptual view of this escape scenario is shown in Figure 12. The figure shows in sketch form the same view as Figure 2. Here, the blue, green, and orange patches represent the main source galaxy, the faint bridge, and the redder interacting companion. The main stellar clumps are drawn in the same positions as seen in that figure. Superimposed is shown in grayscale an illustrative set of contours as they could plausibly look in a case of recent or ongoing tidal stripping by the companion (See lereste2024, Fig. 2, for a real-world observed scenario). The LCE cluster has ionized a surrounding H ii region to a degree where the surrounding neutral Hydrogen, already thin from the tidal interaction, has been completely ionized, allowing LyC photons to well out in a cone-like shape. This cone can, from the results of (almadamonter2024), be up to of the projected surrounding area and still be consistent with the Ly profile reported in (riverathorsen2017).
6 Summary and conclusion
In this work, we present NIRSpec integral field spectroscopy of three pointings at the gravitationally lensed LyC-leaking starburst galaxy, the Sunburst Arc. For each spaxel, we have simultaneously fitted the strongest emission lines to each a single Gaussian profile, all with a shared redshift and FWHM in velocity space, yielding maps of velocity centroid, FWHM, and line fluxes at a pixel resolution of . We constructed maps of dust reddening, ionization parameter , ionization tracer [S II], and shock- and abundance tracer [N II]/H.
We have extracted a stacked spectrum of 5 lensed images of the LyC emitting cluster (the LCE) found within the IFU footprints. We have extracted a magnification-corrected and integrated spectrum of the region within the IFU footprint which covers the largest part of the unlensed galaxy, to mimic a spectrum of the galaxy as it would be observed without gravitational lensing. From these, we were able to compute diagnostic line ratios of interest for LyC escape both for the LCE stack, the full galaxy and in one case also as a 2D map.
We find that:
1.
Line centroid velocity maps show that the Sunburst Arc source galaxy displays a strong, though disturbed, velocity gradient, with an inclination- and lensing-uncorrected velocity span , and a lensing-uncorrected intrinsic velocity dispersion .
2.
The kinematic maps show that the Sunburst Arc is not clearly dominated by neither rotation nor turbulence. The ratio of shearing velocity to intrinsic velocity dispersion is . If the inclination is more than 45°, which is the case for 70% of randomly distributed galaxies, it is dispersion-dominated. In either case, the velocity dispersion is comparatively high for a star forming galaxy, indicating substantial perturbations. We interpret this as a sign of current and/or recent interaction with the companion seen in the upper part of Figure 2, which may have helped spark strong star formation in the main galaxy.
3.
The LCE is clearly visible in maps of ionization indicators such as and . In contrast, it is barely distinguishable in the maps of dust and kinematics. We interpret this as an indication that mechanical feedback is not strong in this region, leaving photoionization as the dominant mechanism clearing away the surrounding H i envelope and facilitating LyC escape.
4.
In a stacked spectrum extracted from five gravitationally lensed images of the LCE cluster, we have identified more than 60 rest-frame optical and red emission lines, spanning from O III 3132 to Pa9 9229, including Balmer lines up to H20 and Paschen lines from Pa9 to Pa22. The stacked spectrum includes a Balmer jump, treated in (welch2025).
5.
The LCE stack has a sulfur deficiency of . Compared to the LzLCS, this places the Sunburst Arc in a range almost exclusively populated by strong LyC emitters. In contrast, the integrated spectrum , a value slightly above the average, in a range which in the LzLCS is populated by both strong and weak leakers and non-leakers (wang2021), showing that contributions from less ionized surrounding regions can dilute this signature beyond detectability even for a strong LyC leaker.
6.
The strong, broad line emission in [O III] and H observed by (mainali2022) does most likely not originate in the LCE itself, but in a Nitrogen-enriched cloud in its vicinity, with a physical distance within a few tens of parsec. This could re-ignite the debates about the role of stellar winds and alternative momentum injection mechanisms in clearing the escape paths for LyC photons.
7.
Based on the extremely low H I column density along the LOS to the LCE clump (sunburst2017), the presence of an interacting companion to the N of the main galaxy(sharon2022), the strong spike in ionization around the LCE and lack of signs of strong mechanical feedback around the LCE cluster; we propose a LyC escape scenario in which tidal interaction with the companion has displaced the majority of the H I envelope around the LCE, leaving the rest to be easily ionized by the strong ionizing emission from the LCE cluster itself.
Appendix A Masks for error estimation and source identification
Here, we include the masks used for noise estimation (Figure 13) and the manual and S/N based mask identifying the observed boundaries of the arc (Figure 14), created as described in Sects. 3.1 and 3.4.2.
Appendix B Emission line identification in the stacked LCE spectrum
Here, we show detailed line identification plots of the stacked spectrum of the Sunburst LCE. We count 60 visually identified lines, plus a number of higher-order Balmer and Paschen lines which have not been labeled here. Figure 15 shows the line ID plot for F100LP/G140H, while Figure 16 shows F170LP/G235H.
In addition, we include in Table 2 a full list of line fluxes measured in the stacked LCE spectrum as described in sect. 3.2. The spectra have been normalized to arbitrary units during stacking, and all fluxes are relative to H.
| Line | F/F(H) | Line | F/F(H) | ||
| He I 3187 | [O I] 6302 | ||||
| [Ne III] 3343 | [S III] 631 | ||||
| O III 3444 | Si II 6347 | ||||
| H18 | [O I] 6365 | ||||
| H17 | [N II] 6550 | ||||
| H16 | H | ||||
| H15 | [N II] 6585 | ||||
| [O II] 3727 | He I 6679 | ||||
| [O II] 3730 | [S II] 6718 | ||||
| H12 | [S II] 6733 | ||||
| H11 | He I | ||||
| H10 | [Ar III] 7138 | ||||
| H9 | He I 7281 | ||||
| [Ne III] | [O II] 7322 | ||||
| H8 | [O II] 7332 | ||||
| [Ne III] 3969 | [Ni II] 7377 | ||||
| H | [Ar III] 7753 | ||||
| He I 4025 | Pa 22 | ||||
| [S II] 4069 | Pa 21 | ||||
| S II] 4076 | Pa 20 | ||||
| H | Pa 19 | ||||
| H | Pa 18 | ||||
| [Fe II] 4360 | O I 8446 | ||||
| [O III] 4363 | Pa 17 | ||||
| He I 4473 | Pa 16 | ||||
| [Fe III] 4658 | Pa 15 | ||||
| [Ar IV] 4713 | Pa 14 | ||||
| [Ar IV] 4741 | Pa 13 | ||||
| H$\star$$\star$footnotemark: | Pa 12 | ||||
| Fe II 4923 | Pa 11 | ||||
| [O III] 4960 | Pa 10 | ||||
| [O III] 5008 | [S III] 9071 | ||||
| He I 5878 | Pa 9 |